I know headquarters wants us to add that new product line, said Dell Havasi, manager of Billings Companys Office ProductsComplete this question by entering your answers in the tabs below. Req 1 to 3 Req 4 Req 5 Req 6A to 6C Req 6D 1. Compute theComplete this question by entering your answers in the tabs below. Req 1 to 3 Req 4 Reg 5 Req 6A to 6C Req 6D 6. Suppose that
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Billing's Company
Ans 1
Particulars Amount ($)
Sales    22,300,000.00 A
Variable Expenses    13,999,600.00 B
Contribution Margin      8,300,400.00 C=A-B
Fixed Expenses      6,115,000.00 D
Operating Income      2,185,400.00 E=C-D
Average Assets      5,575,000.00 F
ROI 39.20% G=E/F
Ans 2
Particulars Amount ($)
Sales      9,650,000.00 H
Variable Expenses      6,272,500.00 I
Contribution Margin      3,377,500.00 J=H-I
Fixed Expenses      2,583,600.00 K
Operating Income         793,900.00 L=J-K
Average Assets      3,857,400.00 M
ROI 20.58% N=L/M
Ans 3
Particulars Amount ($)
Sales    31,950,000.00 O=A+H
Variable Expenses    20,272,100.00 P=B+I
Contribution Margin 11,677,900.00 Q=O-P
Fixed Expenses      8,698,600.00 R=D+K
Operating Income      2,979,300.00 S=Q-R
Average Assets      9,432,400.00 T=F+M
ROI 31.59% U=S/T
Reduction in ROI by 7.61% V=G-U
Ans 4- New product line will decrease ROI by 7.61% so it should not be accepted.
Ans 5- Headquarter is anxious because Office products division is doing very well and adding new product line will benefit company as a whole and also keep the division on ground by reducing its ROI but increasing its operating income in monetary terms.
Ans 6 a
Required Rate of return 14% W
Average Assets      5,575,000.00 See F
Minimum Return         780,500.00 X= W*F
Operating Income      2,185,400.00 See E
Residual Income      1,404,900.00 Y=E-X
Ans 6 b
Required Rate of return 14% See W
Average Assets      3,857,400.00 See M
Minimum Return         540,036.00 Z=W*M
Operating Income         793,900.00 See L
Residual Income         253,864.00 AA= L-Z
Ans 6 c
Required Rate of return 14% See W
Average Assets      9,432,400.00 See T
Minimum Return      1,320,536.00 AB= W*T
Operating Income      2,979,300.00 See S
Residual Income      1,658,764.00 AC=S-AB
Increase in Residual Income by         253,864.00 AD= AC-Y
Ans 6 d- New product line will Increase Residual Income by $ 253,864 so it should be accepted.